OpenStudy (anonymous):

The volume of a cone is three hundred cubic kilometers. If the radius of the base is 6 kilometers, what is the height of the cone? Use three point one four for pi. Show your work and round to the nearest tenth.

OpenStudy (raden):

waw, this cone so big :)

OpenStudy (johnweldon1993):

Hmm, well you know that the volume of a cone can be found using the equation V = (1/3) pi r² h you KNOW V....and r so 300 = (1/3) (3.14) (6²) (h) so rearrange and solve for h :)

OpenStudy (anonymous):

i dont get it

OpenStudy (johnweldon1993):

\[V = \frac{ 1 }{ 3 }\pi r² h\] you know V, pi, and r ...so \[300 = \frac{ 1 }{ 3 } (3.14) (6^2)(h)\] so you would multiply both sides by 3 to get rid of that fraction \[900 = 3.14 (6^2) h\] Divide both sides by (3.14)(6^2) \[\frac{ 900 }{ (3.14)(6^2) } = h\] Do out the math and see that the height of the cone is...?

OpenStudy (johnweldon1993):

Don't forget to round to the nearest tenth :)

OpenStudy (anonymous):

um... its 1.9

OpenStudy (johnweldon1993):

not quite.... 900 ------- 3.14(36) 36 times 3.14 = 113.04 so 900 ---- = ...? 113.04

OpenStudy (anonymous):

7.96

OpenStudy (johnweldon1993):

perfect...and rounded to the nearest tenth would be...

OpenStudy (anonymous):

8

OpenStudy (johnweldon1993):

And you're all set! :)

OpenStudy (anonymous):

thank you so much

OpenStudy (johnweldon1993):

anytime! :)